OpenStudy (anonymous):

Solve for m. Then find the value of m for each value of n. m + 3n = 7; n = -2, 0, 1

OpenStudy (austinl):

We have, \(m+3n=7\) To solve for m, we would simply subtract \(3n\) from both sides. \(m=-3n+7\) Then we would plug in the values for \(n\) of -2,0,1. \(m=-3(-2)+7=~?\) \(m=-3(0)+7=~?\) \(m=-3(1)+7=~?\)

OpenStudy (austinl):

Does this make sense @ElmerHigglesworth ?

OpenStudy (anonymous):

I'm sure it does, but my primitive 8th grader brain is losing power.

OpenStudy (austinl):

What does the first equation of \(m\) equal? \(m=-3(-2)+7\)